TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-18 22:45:00
subject: 3\01 Pt-1 ISS On Orbit Status 01-03-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
3\01 ISS On-Orbit Status 01-03-2003
Part 1 of 2

ISS On-Orbit Status 3/1/03

All ISS systems continue to function nominally, except as noted 
previously or below. Space Day 98 for Expedition 6.

A regular off-duty Saturday for the crew, except for weekend routine 
tasks, necessary repair work and daily maintenance.

After wake-up (1:00am EST), morning inspection, morning hygiene and
breakfast, the ISS residents completed the weekly 3-hr. house 
cleaning, designed to keep the crew's orbital outpost spic and span. 
[The "uborka stantsii" features removal of food waste products, 
cleaning of compartments with vacuum cleaner, wet cleaning of surfaces 
with disinfectants and cleaning of fan screens to avoid temperature 
rises.]

FE-1 Nikolai Budarin terminated the regeneration process for 
absorption bed #2 of the BMP harmful impurities filtration unit. Both 
filter channels are now back in Purify mode.

Budarin also conducted the weekly inspection of the BRPK 
air/condensate separator of the SRVK water processing system. [The 
BRPK separator contains porous Cermet hydrophilic (water attracting) 
tubes through which the gas-liquid mixture from the heat exchanger 
moves. They separate the air from the condensate, but when the 
separator exceeds its service life or is malfunctioning, incomplete 
separation of the atmospheric condensate occurs, and the water then 
collects under the "sheet" of porous fluoroplastic. This is the main 
focus of the regular inspection.]

FE-2/SO Don Pettit performed the daily routine maintenance of the SOSh 
life support system in the Service Module (SM) and prepared the IMS 
(inventory management system) auto import/export files.

All crewmembers performed their regular daily physical exercise on 
TVIS (treadmill with vibration isolation and stabilization), RED 
(resistive exercise device), CEVIS (cycle ergometer with vibration 
isolation) and VELO bike, with load trainer.

Yesterday (9:15am EST) the crew participated in two outstanding 
live-TV interviews, the first with USA Today (Traci Watson), the other 
with KPTV-TV, Portland, OR (Pete Ferryman).

The cause of the off-nominal shutdown of the Elektron, Vozdukh and 
SKV-1 on 2/27 was traced by MCC-M and RSC-Energia to the primary data 
bank (BD) of the BITS2-12 onboard measurement telemetry system. BD-SU 
mode was power-cycled (turned off and on), and BITS was also switched 
to the backup data bank. This reportedly corrected the problems with 
the BITS system.  Vozdukh, Elektron and SKV-1 were running again by 
yesterday noon. Analysis of the event is continuing.

The Lab CDRA (CO2 removal assembly) is off. To configure it for 
extended deactivation, the crew today disconnected the ITCS LTL 
(internal thermal control system/low temperature loop).

The uplink of the GNC (guidance, navigation & control) software patch 
to correct the "RGA data staleness" problem with RGA (rate gyro 
assembly) data processing in the new R3 software was deferred to 
today, due to ground testing issues last night. [Also uplinked were 
contingency procedures for restoring communications with MCC-H for the 
event of an unexpected failure of the primary GNC MDM during the time 
in which the backup GNC MDM is being configured for patch loading. In 
this situation, there would be no GNC MDMs, a loss of pointing data 
and, subsequently,a loss of S-band comm. The contingency procedures 
would guide the crew in configuring for low rate S-band comm in this 
case.]

Today's occurrence of the Bradfield Meteor Shower will be visible from 
any earth-facing window of the ISS. The appropriate viewing times 
(4:39 to 4:58pm EST) were uplinked in case the crew wants to observe 
the show,- close to their sleep time. At that time, the Station will 
be over the southern coast of Africa and Madagascar.

Science Update (Expedition Six -- 13th):
The crew was congratulated by the Lead Increment Scientist on 
verifying the modification necessary to ensure recording of the FOOT 
data. They were thanked for completing one more PuFF run, downloading 
EVARM data, troubleshooting the MSG rack, locating the ProMISS 
(protein crystal growth monitoring by digital holography) hardware, 
and taking and downlinking CEO and Renal photos. Work is underway on 
relaxing the six-hour cool down time after MSG deactivation. This will 
allow the remaining testing to proceed more rapidly.

Extra-Vehicular Activity Radiation Monitors (EVARM): The downlinked 
badge read data continues to give invaluable insight into ISS.

GASMAP/Pulmonary Function in Flight (PuFF): The PuFF team sent "gold 
stars" to Don Pettit for his performance in lung volume calibration. 
The downlink video was also appreciated.

Renal (Kidney) Stone Experiment: Completed for Increment 6. The Renal 
Stone Experiment Team thanked the crew for taking photos of the Renal 
Stone hardware items. Bowersox and Pettit continue taking their 
stone-prevention (or placebo) pills.

Human Research Facility/Workstation (HRF WS): Continuing.

Foot/Ground Reaction Forces During Space Flight (FOOT): The FOOT team
thanked Bowersox for his much-needed help troubleshooting the ADAS and 
FOOT software. A work-around in the nominal procedure will be added 
for avoiding the data loss during future data collection sessions.

Pore Formation and Mobility Investigation (PFMI): On hold until MSG
(microgravity science glovebox) is operational. Remaining in MSG until
further notice.

Space Acceleration Measurement System (SAMS): SAMS continues to 
collect acceleration data for vibratory characterization of the 
microgravity environment, including the recent reboost activities.

Microgravity Acceleration Measurement System (MAMS): MAMS continues
measurement of microgravity environment in the quasi-steady regime for
general characterization.

Protein Crystal Growth-Single Locker Thermal Enclosure System 
(PCG-STES):  Temperatures are nominal.

Investigating the Structure of Paramagnetic Aggregates from Colloidal
Emulsions (InSPACE): On hold until MSG is operational.

Materials ISS Experiment (MISSE): In progress. Deployed outside. 
Nominal and collecting data.

Zeolite Crystal Growth (ZCG): ZCG has finished science operations for 
Inc 6.

EarthKAM (EK): Planning additional operations cycles for Increment 6.

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.